# Estimating the Probabilities of Rare Outputs in Language Models

### v1

@@ 3.2.1 Quadratic Logit Decomposition (QLD)
Choice of direction.  We rely on the following two assumptions for QLD to perform well: 1) ${\mathbf{a}}$ and ${\mathbf{b}}$ are independent, and 2) the contribution towards the output behavior is split roughly equally between these two terms (see Appendix C for more discussion). After some initial experimentation with a variety of candidate directions, [^4] Other candidate directions included 1) the $t$-th column of ${\bm{W}}_{U}$ pulled back into whitened space and 2) the expectation of $\mathcal{N}(0,\mathrm{Id}_{d})$ conditioned on lying in the whitened acceptance region. we decided to set ${\bm{d}}$ to be the direction of the shortest vector in whitened space that results in the model outputting $t$. It can also be thought of as the maximum likelihood value of ${\mathbf{v}}$ under a Gaussian prior, conditioned on observing the model output $t$. Appendix D describes the algorithm we use to compute ${\bm{d}}$.
@@ 4 Experimental Setup
For each of the $8$ distributions (listed in Table 1) and for each model, we generate ground-truth token probabilities by running forward passes on $2^{32}$ random samples. We then select a random set of 256 tokens among those with ground-truth probabilities between $10^{-9}$ and $10^{-5}$, and we test all of our methods on these tokens. [^5] The hex distribution only had $159$ and $135$ such tokens for the $2$- and $4$-layer models, respectively, so we instead used every token that fell into the given range. See Appendix E for more details.
@@ 4.2 Affine Fits
Many methods often report estimates of 00, but $D_{\mathrm{IS}}$ is undefined for $q=0$. To address this, we fit a transformation $x\mapsto ax^{c}+b$ to the outputs of each method, where $a,b$ and $c$ are chosen to minimize Itakura–Saito loss. $ax^{c}$ can be thought of an affine transformation in log-space, and adding $b$ prevents values from being too small while barely affecting larger outputs. To ensure that this transformation is not overfitting to the particular set of 256 tokens, we report the leave-one-out cross-validation (LOOCV) loss of each method. We train a separate fit for each (method, input distribution) pair.
Note that the Gaussian Logit Difference method has a special functional form of its fit ($(\mu,\sigma)\mapsto\exp\left(-\left(\frac{a\mu}{\sigma+\epsilon}\right)^{2}+b\right)+c$ instead of $x\mapsto ax^{c}+b$) but is otherwise evaluated in the same way.
@@ 6.4 Limitations
There are two main limitations of our experimental setup. First, we only use input distributions that factor into independent tokens. This choice is necessary for the definition of ITGIS. It is also very convenient for the implementation of MHIS, because it gives efficient sampling access to the proposal distribution. To move beyond independent token input distributions, we could define the input distribution to be the output of a separate generative model and adapt some of the current estimation methods appropriately (GLD and QLD are input-agnostic and could be applied in their current form).
Second, we only study model behaviors that consist of a single token sampled at temperature 00. This is unrealistic because in practice, if we were concerned about specific single-token outputs, it would be easy to block them with a filter. In contrast, the types of behaviors we actually worry about likely involve long chains of autoregressive generation or interaction with the external world (e.g., when forming and executing a plan). We are excited to see future work extending our setting in this direction.
@@ 7 Related Work
Finally, there is a large body of work applying adversarial training to improve worst-case model performance (Bai et al., 2021; Goodfellow et al., 2014; Ilyas et al., 2019), especially in the context of language models (Madry, 2017; Liu et al., 2020). Perez et al. (2022) explores using language models themselves to aid in red-teaming other models. Latent adversarial training (Casper et al., 2024) generalizes standard adversarial training by optimizing over perturbations in activation space; this means that, like activation extrapolation methods, it can be effective even when the adversarial training search problem over input space is hard.
@@ 9 Acknowledgements
Paul Christiano played a significant role in advising the project. We are grateful for intermediate theoretical contributions from David Matolsci and George Robinson. Thanks additionally to Jean-Stanislas Denain and Eric Neyman for feedback on a draft.
@@ Appendix A Input distributions
icl 29 A simple in-context learning prompt of the form ‘A for _ R for _ C for _Y for _ ? for’, where the underscores are replaced with random tokens that start with the corresponding letter (weighted by frequency in English text), and the ? is replaced with a uniformly random letter.
@@ Appendix B Full Estimation Algorithms
    Draw a batch of samples from from $q^{(j-1)}$. Use the empirical mean of their gradient to estimate $\hat{s}_{i}^{(j)}(x_{i})=\operatorname{\mathbb{E}}_{{\mathbf{x}}^{\prime}\sim q^{(j-1)}}[\nabla_{{\mathbf{x}}^{\prime}}M_{t}({\mathbf{x}}^{\prime})]_{i,x_{i}}$ for all $i\in[k],x_{i}\in\mathcal{V}$.
@@
The final output of the method is the average value of this importance sampling estimator across all batches. The number of batches and samples per batch is specified in Appendix F.
@@
    $p_{k}(x^{\prime}_{i})\cdot\exp\left(\frac{{\mathbf{g}}_{x^{\prime}_{i}}}{T}\right).$
@@
The right-hand side can be estimated using the $n$ samples we already have from (approximately) $q$. In practice, the way we estimate the normalizing constant does not matter much (most of the error comes from other steps).
@@
This proportion can be computed in $\tilde{O}(n)$ time—we don’t need to explicitly iterate over all $n^{2}$ pairs. By the convexity of the acceptance region $S$, for any fixed ${\mathbf{b}}$ there is a single interval $[\ell,r]$ such that $a\in[\ell,r]\Leftrightarrow a{\bm{d}}+{\mathbf{b}}\in S$. We can efficiently compute the bounds of this interval for every sample ${\mathbf{b}}^{(j)}$ by solving a linear system of inequalities, and then we can calculate how many ${\mathbf{a}}^{(i)}$ fall into each range in $O(\log n)$ time after sorting. Thus, the computational cost of QLD is dominated by running $n$ forwards passes of $M$ to generate the samples ${\mathbf{u}}^{(1)},\dots,{\mathbf{u}}^{(n)}$.
@@ Appendix D Computing the Shortest Accepting Vector
2:  for $\mathrm{step\_cnt}\in\{1,\dots,100\cdot n_{\text{reps}}\}$ do
@@ Appendix E Ground Truth Token Distribution
Figure 5: The ground truth probabilities of tokens for each distribution and model size, sorted from most to least probable (the height of the curve at position $x$ is the probability of the $x$-th most common token). Any tokens that appeared 00 times across all $2^{32}$ samples are not plotted.

### v2

@@ 3.2.1 Quadratic Logit Decomposition (QLD)
Choice of direction.  We rely on the following two assumptions for QLD to perform well: 1) ${\mathbf{a}}$ and ${\mathbf{b}}$ are independent (so that our estimate is unbiased), and 2) the contribution towards the output behavior is split roughly equally between these two terms (to minimize the variance of our estimate). See Appendix C for more discussion of this motivation. After some initial experimentation with a variety of candidate directions, [^4] Other candidate directions included 1) the $t$-th column of ${\bm{W}}_{U}$ pulled back into whitened space and 2) the expectation of $\mathcal{N}(0,\mathrm{Id}_{d})$ conditioned on lying in the whitened acceptance region. we decided to set ${\bm{d}}$ to be the direction of the shortest vector in whitened space that results in the model outputting $t$. It can also be thought of as the maximum likelihood value of ${\mathbf{v}}$ under a Gaussian prior, conditioned on observing the model output $t$. Appendix D describes the algorithm we use to compute ${\bm{d}}$.
@@ 4 Experimental Setup
For each of the $8$ distributions (listed in Table 1) and for each model, we generate ground-truth token probabilities by running forward passes on $2^{32}$ random samples. We then select a random set of 256 tokens among those with ground-truth probabilities between $10^{-9}$ and $10^{-5}$, and we test all of our methods on these tokens.
@@ 4.2 Affine Fits
Many methods often report estimates of 00, but $D_{\mathrm{IS}}$ is undefined for $q=0$. To address this, we fit a transformation $x\mapsto ax^{c}+b$ to the outputs of each method, where $a,b$ and $c$ are chosen to minimize Itakura–Saito loss. $ax^{c}$ can be thought of an affine transformation in log-space, and adding $b$ prevents values from being too small while barely affecting larger outputs. To ensure that this transformation is not overfitting to the particular set of 256 tokens, we report the leave-one-out cross-validation (LOOCV) loss of each method. We train a separate fit for each (method, input distribution) pair.  [^5] Note that the Gaussian Logit Difference method has a special functional form of its fit ($(\mu,\sigma)\mapsto\exp\left(-\left(a\mu/(\sigma+\epsilon)\right)^{2}+b\right)+c$ instead of $x\mapsto ax^{c}+b$) but is otherwise evaluated in the same way.
@@ 6.4 Limitations
There are two main limitations of our experimental setup. First, we only use input distributions that factor into independent tokens. This choice is necessary for the definition of ITGIS. It is also very convenient for the implementation of MHIS, because it gives efficient sampling access to the proposal distribution. To move beyond independent token input distributions, we could define the input distribution to be the output of a separate generative model and adapt some of the current estimation methods appropriately.
Second, we only study model behaviors that consist of a single token sampled at temperature 00. This is unrealistic because in practice, if we were concerned about specific single-token outputs, it would be easy to filter them out. In contrast, the types of behaviors we actually worry about likely involve long chains of autoregressive generation or interaction with the external world (e.g., when forming and executing a plan). We are excited to see future work extending our setting in this direction.
@@ 7 Related Work
Finally, there is a large body of work applying adversarial training to improve worst-case model performance (Bai et al., 2021; Goodfellow et al., 2014; Ilyas et al., 2019), especially in the context of language models (Madry, 2017; Liu et al., 2020). Perez et al. (2022) explores using language models themselves to aid in red-teaming other models. Latent adversarial training (Casper et al., 2024; Sheshadri et al., 2024) generalizes standard adversarial training by optimizing over perturbations in activation space; this means that, like activation extrapolation methods, it can be effective even when the adversarial training search problem over input space is hard.
@@ 9 Acknowledgements
Paul Christiano played a significant role in advising the project. We are grateful for intermediate theoretical contributions from David Matolcsi and George Robinson. Thanks additionally to Jean-Stanislas Denain and Eric Neyman for feedback on a draft.
@@ Appendix A Input distributions
icl 29 A simple in-context learning prompt of the form ‘A for _ R for _ C for _Y for _ ? for’, where the underscores are replaced with random tokens that start with the corresponding letter (weighted by frequency in English text), and the ? is replaced with a uniformly random letter. The letters spell out ‘ARCTHEORY’.
@@ Appendix B Full Estimation Algorithms
    Draw a batch of samples from from $q^{(j-1)}$. Use the empirical mean of their gradient to estimate $\hat{s}_{i}^{(j)}(x)=\operatorname{\mathbb{E}}_{{\mathbf{x}}^{\prime}\sim q^{(j-1)}}[\nabla_{{\mathbf{x}}^{\prime}}M_{t}({\mathbf{x}}^{\prime})]_{i,x}$ for all $i\in[k],x\in\mathcal{V}$.
@@
The final output of the method is the average value of this importance sampling estimator across all batches. The number of batches and samples per batch is specified in Appendix F. See Algorithm 1 for pseudocode.
Algorithm 1 Independent Token Gradient Importance Sampling (ITGIS)
0:  Model $M$, target token $t$, input length $k$, token distributions $p_{1},\dots,p_{k}$, temperature $T$, iterations $n$, batch size $B$
1:  $s^{(0)}_{i}\leftarrow\mathbf{0}\in\mathbb{R}^{|\mathcal{V}|}$ for all $i\in[k]$ Initialize score functions
2:  $\mathrm{estimates}\leftarrow[]$
2:
3:  for $j\leftarrow 1$ to $n$ do
4:     for $i\leftarrow 1$ to $k$ do
5:        $q^{(j-1)}_{i}(x)\leftarrow p_{i}(x)\cdot\exp(s^{(j-1)}_{i}(x)/T)$ for all $x\in\mathcal{V}$
6:        Normalize $q^{(j-1)}_{i}$ to have sum $1$
7:     end for
7:
8:     Sample $B$ inputs $\{{\mathbf{x}}^{(b)}\}_{b=1}^{B}$ from $q^{(j-1)}_{1}\times\dots\times q^{(j-1)}_{k}$
9:     for $i\leftarrow 1$ to $k$ do
10:        $\hat{s}^{(j)}_{i}(x)\leftarrow\frac{1}{B}\sum_{b=1}^{B}[\nabla_{{\mathbf{x}}}M_{t}({\mathbf{x}}^{(b)})]_{i,x}$ for all $x\in\mathcal{V}$
11:     end for
11:
12:     $\alpha\leftarrow 0.9$
13:     for $i\leftarrow 1$ to $k$ do
14:        $s^{(j)}_{i}\leftarrow\frac{\hat{s}^{(j)}_{i}+\alpha\hat{s}^{(j-1)}_{i}+\dots+\alpha^{j-1}\hat{s}^{(1)}_{i}}{1+\alpha+\alpha^{2}+\dots+\alpha^{j-1}}$
15:     end for
15:
16:     $\mathrm{estimate}\leftarrow\frac{1}{B}\sum_{b=1}^{B}\frac{\prod_{i=1}^{k}p_{i}({\mathbf{x}}^{(b)}_{i})}{\prod_{i=1}^{k}q^{(j-1)}_{i}({\mathbf{x}}^{(b)}_{i})}\mathbbm{1}[M({\mathbf{x}}^{(b)})=t]$
17:     Append $\mathrm{estimate}$ to $\mathrm{estimates}$
18:  end for
18:
19:  return  $\frac{1}{n}\sum_{j=1}^{n}\mathrm{estimates}[j]$
@@
    $p_{i}(x^{\prime}_{i})\cdot\exp\left(\frac{{\mathbf{g}}_{x^{\prime}_{i}}}{T}\right).$
@@
The right-hand side can be estimated using the $n$ samples we already have from (approximately) $q$. In practice, the way we estimate the normalizing constant does not matter much (most of the error comes from other steps). See Algorithm 2 for pseudocode.
Algorithm 2 Metropolis–Hastings Importance Sampling (MHIS)
0:  Model $M$, target token $t$, input length $k$, token distributions $p_{1},\dots,p_{k}$, temperature $T$, number of burn-in steps $n_{\text{burn}}$, number of samples $n$
1:  Initialize ${\mathbf{x}}$ by sampling from $p_{1}\times\dots\times p_{k}$
2:  $\mathrm{samples}\leftarrow[]$
3:  for $\mathrm{step}\leftarrow 1$ to $n_{\text{burn}}+n$ do
4:     $i\leftarrow\mathrm{Unif}([k])$ Choose random position
5:     ${\mathbf{g}}\leftarrow[\nabla_{{\mathbf{x}}}M_{t}({\mathbf{x}})]_{i}$
6:     Sample ${\textnormal{x}}^{\prime}_{i}$ from $\propto p_{i}(x^{\prime}_{i})\cdot\exp({\mathbf{g}}_{x^{\prime}_{i}}/T)$ Proposed token
7:     ${\mathbf{x}}^{\prime}\leftarrow({\mathbf{x}}_{1},\dots,{\textnormal{x}}^{\prime}_{i},\dots,{\mathbf{x}}_{k})$ Proposed new state
7:      # Compute acceptance ratio $r=\frac{q({\mathbf{x}}^{\prime})\phi({\mathbf{x}}|{\mathbf{x}}^{\prime})}{q({\mathbf{x}})\phi({\mathbf{x}}^{\prime}|{\mathbf{x}})}$
8:     $r\leftarrow\mathrm{AcceptanceRatio}(M,t,T,(p_{1},\dots,p_{k}),{\mathbf{x}},{\mathbf{x}}^{\prime})$
9:     if $\mathrm{Unif}(0,1)<r$ then
10:        ${\mathbf{x}}\leftarrow{\mathbf{x}}^{\prime}$ Accept proposal
11:     end if
11:
12:     if $\mathrm{step}>n_{\text{burn}}$ then
13:        Append ${\mathbf{x}}$ to $\mathrm{samples}$
14:     end if
15:  end for
15:
16:  $Z\leftarrow\left(\frac{1}{n}\sum_{j=1}^{n}\exp(-M_{t}(\mathrm{samples}[j])/T)\right)^{-1}$ Est. normalizing constant
17:  return  $\frac{1}{n}\sum_{j=1}^{n}\frac{\exp(M_{t}(\mathrm{samples}[j])/T)}{Z}\mathbbm{1}[M(\mathrm{samples}[j])=t]$
@@
This proportion can be computed in $\tilde{O}(n)$ time—we don’t need to explicitly iterate over all $n^{2}$ pairs. By the convexity of the acceptance region $S$, for any fixed ${\mathbf{b}}$ there is a single interval $[\ell,r]$ such that $a\in[\ell,r]\Leftrightarrow a{\bm{d}}+{\mathbf{b}}\in S$. We can efficiently compute the bounds of this interval for every sample ${\mathbf{b}}^{(j)}$ by solving a linear system of inequalities, and then we can calculate how many ${\mathbf{a}}^{(i)}$ fall into each range in $O(\log n)$ time after sorting. Thus, the computational cost of QLD is dominated by running $n$ forwards passes of $M$ to generate the samples ${\mathbf{u}}^{(1)},\dots,{\mathbf{u}}^{(n)}$. See Algorithm 3 for pseudocode.
Algorithm 3 Quadratic Logit Decomposition (QLD)
0:  Model $M$ with unembed matrix ${\bm{W}}_{U}$, target token $t$, sample size $n$
1:  Sample $n$ inputs ${\mathbf{x}}^{(1)},\dots,{\mathbf{x}}^{(n)}$ from input distribution
2:  Compute pre-unembed activations ${\mathbf{v}}^{(i)}\in\mathbb{R}^{d}$ by running $M$ on ${\mathbf{x}}^{(i)}$ for $i\in[n]$
3:  ${\bm{\mu}}\leftarrow\frac{1}{n}\sum_{i=1}^{n}{\mathbf{v}}^{(i)}$
4:  ${\bm{\Sigma}}\leftarrow\frac{1}{n}\sum_{i=1}^{n}({\mathbf{v}}^{(i)}-{\bm{\mu}})({\mathbf{v}}^{(i)}-{\bm{\mu}})^{\top}$
5:  Find ${\bm{A}}$ such that ${\bm{A}}{\bm{A}}^{\top}={\bm{\Sigma}}+\epsilon\cdot\mathrm{Id}_{d}$ via Cholesky decomposition
6:  for $i\leftarrow 1$ to $n$ do
7:     ${\mathbf{u}}^{(i)}\leftarrow{\bm{A}}^{-1}({\mathbf{v}}^{(i)}-{\bm{\mu}})$ Whitened activations
8:  end for
8:
9:  ${\bm{d}}\leftarrow\mathrm{ShortestAcceptingVector}({\bm{A}},{\bm{\mu}},{\bm{W}}_{U},t)$ See Appendix D
10:  for $i\leftarrow 1$ to $n$ do
11:     ${\mathbf{a}}^{(i)}\leftarrow{\bm{d}}{\bm{d}}^{\top}{\mathbf{u}}^{(i)}$ Parallel component
12:     ${\mathbf{b}}^{(i)}\leftarrow{\mathbf{u}}^{(i)}-{\mathbf{a}}^{(i)}$ Perpendicular component
13:  end for
13:
14:  Sort $\{{\mathbf{a}}^{(i)}\}_{i=1}^{n}$ in ascending order
15:  $\mathrm{count}\leftarrow 0$
16:  for $j\leftarrow 1$ to $n$ do
16:      # Solve linear inequalities to get $[\ell_{j},r_{j}]$ such that $a\in[\ell_{j},r_{j}]\Leftrightarrow a{\bm{d}}+{\mathbf{b}}^{(j)}\in S$
17:     $[\ell_{j},r_{j}]\leftarrow\mathrm{FindAcceptanceInterval}({\mathbf{b}}^{(j)},{\bm{d}},{\bm{A}},{\bm{\mu}},{\bm{W}}_{U},t)$
18:     $\mathrm{count}\leftarrow\mathrm{count}+|\{i:{\mathbf{a}}^{(i)}\in[\ell_{j},r_{j}]\}|$ Binary search for bounds
19:  end for
19:
20:  return  $\mathrm{count}/n^{2}$
@@ Appendix C Principles for choosing a decomposition in QLD
When we run QLD with many more samples than its standard budget of $2^{16}$, its performance improves but plateaus at a level that is still worse than the sampling methods. This shows that QLD is currently limited by a poor independence assumption, as the error arising from an unequal split of contribution should vanish with a sufficient number of samples.
@@ Appendix D Computing the Shortest Accepting Vector
0:  Half-spaces $H_{1},\dots,H_{t-1},H_{t+1},\dots,H_{|\mathcal{V}|}$, number of repetitions $n_{\text{reps}}$
@@
2:  for $\mathrm{step\_cnt}\leftarrow 1$ to $100\cdot n_{\text{reps}}$ do
@@ Appendix E Ground Truth Token Distribution
Figure 5: The ground truth probabilities of tokens for each distribution and model size, sorted from most to least probable (the height of the curve at position $x$ is the probability of the $x$-th most common token). Any tokens that appeared 00 times across all $2^{32}$ samples are not plotted. The hex distribution only had $159$ and $135$ tokens in the range $[10^{-9},10^{-5}]$ for the $2$- and $4$-layer models, respectively, so we used every such token instead of sampling $256$ of them.